# Large Language Models for Secure Code Assessment: A Multi-Language Empirical Study

### v1

@@ II-C Large Language Models
GPT-3.5 Turbo 0.62 0.59 0.77 0.67 0.57 0.55 0.82 0.66 0.72 0.73 0.70 0.71 0.61 0.58 0.77 0.66
GPT-4 Turbo 0.81 0.81 0.82 0.81 0.79 0.73 0.92 0.81 0.79 0.75 0.87 0.81 0.78 0.71 0.93 0.81
GPT-4o 0.80 0.72 0.97 0.83 0.71 0.64 0.99 0.78 0.62 0.57 1.00 0.72 0.63 0.57 0.99 0.73
CodeLlama-7b 0.62 0.58 0.89 0.70 0.57 0.54 0.93 0.68 0.54 0.52 0.98 0.68 0.57 0.55 0.84 0.66
CodeLlama-13b 0.60 0.76 0.28 0.41 0.66 0.67 0.65 0.66 0.65 0.62 0.77 0.69 0.66 0.62 0.81 0.71
Gemini 1.5 Pro 0.80 0.79 0.83 0.81 0.69 0.65 0.84 0.73 0.75 0.71 0.83 0.77 0.68 0.62 0.97 0.75
@@
GPT-3.5 Turbo 0.63 0.59 0.87 0.70 0.57 0.54 0.89 0.67 0.74 0.75 0.71 0.73 0.61 0.57 0.87 0.69
GPT-4 Turbo 0.75 0.73 0.79 0.76 0.78 0.72 0.89 0.80 0.78 0.73 0.87 0.80 0.71 0.66 0.87 0.75
GPT-4o 0.76 0.70 0.92 0.80 0.66 0.60 0.95 0.73 0.58 0.54 1.00 0.70 0.59 0.55 0.97 0.70
CodeLlama-7b 0.70 0.67 0.76 0.72 0.58 0.56 0.79 0.65 0.54 0.52 0.97 0.68 0.57 0.55 0.76 0.64
CodeLlama-13b 0.57 0.69 0.24 0.35 0.59 0.59 0.61 0.60 0.59 0.59 0.63 0.61 0.54 0.53 0.63 0.58
Gemini 1.5 Pro 0.76 0.79 0.71 0.75 0.61 0.59 0.71 0.64 0.74 0.70 0.84 0.76 0.66 0.60 0.95 0.73
@@
GPT-3.5 Turbo 0.56 0.55 0.68 0.61 0.47 0.48 0.70 0.57 0.69 0.68 0.73 0.70 0.57 0.56 0.68 0.61
GPT-4 Turbo 0.83 0.78 0.91 0.84 0.75 0.67 0.98 0.80 0.80 0.75 0.89 0.81 0.78 0.70 0.98 0.82
GPT-4o 0.83 0.75 1.00 0.85 0.70 0.63 1.00 0.77 0.64 0.58 1.00 0.73 0.66 0.59 1.00 0.75
CodeLlama-7b 0.59 0.55 0.93 0.69 0.59 0.55 0.98 0.70 0.52 0.51 0.95 0.67 0.52 0.51 0.86 0.64
CodeLlama-13b 0.65 0.93 0.32 0.47 0.69 0.67 0.77 0.72 0.65 0.60 0.86 0.71 0.77 0.72 0.89 0.80
Gemini 1.5 Pro 0.77 0.73 0.86 0.79 0.70 0.67 0.80 0.73 0.78 0.79 0.77 0.78 0.73 0.65 1.00 0.79
@@
GPT-3.5 Turbo 0.60 0.57 0.75 0.65 0.60 0.56 0.86 0.68 0.72 0.71 0.75 0.73 0.61 0.59 0.75 0.66
GPT-4 Turbo 0.85 0.82 0.89 0.85 0.83 0.76 0.97 0.85 0.78 0.71 0.94 0.81 0.81 0.73 0.97 0.83
GPT-4o 0.86 0.78 1.00 0.88 0.75 0.67 1.00 0.80 0.62 0.57 1.00 0.73 0.60 0.55 1.00 0.71
CodeLlama-7b 0.61 0.57 0.92 0.70 0.57 0.54 0.94 0.69 0.57 0.54 1.00 0.70 0.60 0.56 0.94 0.70
CodeLlama-13b 0.61 0.83 0.28 0.42 0.62 0.63 0.61 0.62 0.68 0.64 0.83 0.72 0.69 0.64 0.89 0.74
Gemini 1.5 Pro 0.88 0.85 0.92 0.88 0.72 0.67 0.89 0.76 0.72 0.67 0.89 0.76 0.68 0.61 0.97 0.75
@@
GPT-3.5 Turbo 0.74 0.69 0.86 0.77 0.69 0.63 0.93 0.75 0.81 0.85 0.76 0.80 0.72 0.67 0.90 0.76
GPT-4 Turbo 0.84 0.88 0.79 0.84 0.83 0.77 0.93 0.84 0.78 0.74 0.86 0.79 0.72 0.67 0.90 0.76
GPT-4o 0.67 0.61 0.97 0.75 0.66 0.59 1.00 0.74 0.53 0.52 1.00 0.68 0.53 0.52 0.97 0.67
CodeLlama-7b 0.64 0.58 0.97 0.73 0.55 0.53 0.97 0.68 0.52 0.51 1.00 0.67 0.66 0.60 0.90 0.72
CodeLlama-13b 0.53 0.57 0.28 0.37 0.69 0.72 0.62 0.67 0.66 0.64 0.72 0.68 0.62 0.58 0.86 0.69
Gemini 1.5 Pro 0.88 0.87 0.90 0.88 0.64 0.60 0.86 0.70 0.71 0.67 0.83 0.74 0.60 0.56 1.00 0.72
@@
GPT-3.5 Turbo 0.60 0.57 0.74 0.65 0.58 0.56 0.76 0.65 0.65 0.69 0.57 0.62 0.58 0.57 0.69 0.62
GPT-4 Turbo 0.80 0.86 0.71 0.78 0.79 0.76 0.83 0.80 0.82 0.83 0.81 0.82 0.83 0.78 0.93 0.85
GPT-4o 0.82 0.75 0.98 0.85 0.79 0.70 1.00 0.82 0.68 0.61 1.00 0.76 0.71 0.64 1.00 0.78
CodeLlama-7b 0.56 0.54 0.88 0.67 0.54 0.52 0.95 0.67 0.54 0.52 1.00 0.68 0.56 0.54 0.76 0.63
CodeLlama-13b 0.60 0.75 0.29 0.41 0.71 0.78 0.60 0.68 0.68 0.65 0.79 0.71 0.65 0.62 0.79 0.69
Gemini 1.5 Pro 0.75 0.74 0.76 0.75 0.76 0.70 0.93 0.80 0.77 0.74 0.83 0.79 0.71 0.65 0.95 0.77
II-D Classification Methods
Our study examined three classification approaches:
  • Vulnerability Detection: A simple binary classification method in which the LLM is tasked to categorize an input snippet into either “vulnerable” or “not vulnerable”.
  • CWE classification (zero-shot): A multi-class classification method in which the LLM is tasked to categorize an input snippet as pertaining to a specific CWE. The zero-shot setting provides only a task description in natural language as part of the input prompt.
  • CWE classification (few-shot): Same multi-class classification method as before, but using a few-shot setting in which the LLM is provided with code examples from the list of the top 25 CWEs [25].
II-E Prompt and Input Data
@@
In our study, the final prompts combine one system prompt with one user prompt. Table III reports all the prompts that were used for our experiments. We considered two system prompts: VD-Sys1 gives the LLM a detailed role for vulnerability detection. In contrast, VD-Sys2 had a more generic role, as described by Zhou et al. [37]. Similarly, we designed two different types of user prompts to detect a vulnerability: VD-User1 presents a concise request for vulnerability detection, whereas VD-User2 provides more descriptive instruction as used by Zhou et al. [37].
For the classification of CWE (RQ2), we introduced two system prompts (CWE-Sys1 and CWE-Sys2) and two user prompts (CWE-UserZ and CWE-UserF).
CWE-Sys1 is a similar prompt compared to VD-Sys1, which provides a more specific role for CWE classification. Conversely, CWE-Sys2 is the same prompt as VD-Sys2 [37].
CWE-UserZ is a user prompt for CWE classification with a zero-shot setting (i.e., without any examples). CWE-UserF is for CWE classification experiments with the few-shot setting, and it has a list of the top 25 CWEs comparable to CWE-UserZ. To form the final prompt text, we concatenated an input code snippet with one of our described prompt strategies.
II-F Procedure and Metrics
To answer RQ1 and RQ2, we executed each LLM on all code snippets of our datasets, using the prompt strategies described in Section II-E. As evaluation metrics, we employed four metrics, namely accuracy, recall, precision, and $F_{1}$ score, commonly used to evaluate the performance of LLMs in vulnerability detection.
In our evaluation, valid detections/classifications were treated as the positive class, while detection/classification augmentations were considered the negative class. We accounted for CWE hierarchical relationships, considering positive scores for cases where the LLM predicted a code snippet labeled with a specific CWE (e.g., CWE-77) as one of its child CWEs (e.g., CWE-78), according to the MITRE CWE database [38]. For CWE classification (RQ2), we evaluated the overall performance by macro-averaging the precision parameter across all classes in the precision, recall, and $F_{1}$ score.
Overall, in our evaluation, we computed 18,144 predictions (6 LLMs $\times$ 378 code snippets $\times$ 8 prompt configurations), approximately 10 hours of computing time.
Table V: RQ2: Effectiveness comparison of six LLMs using four prompt configurations for CWE classification. The best scores are indicated in bold.
LLM / Prompt CWE-Sys1 + CWE-UserZ CWE-Sys2 + CWE-UserZ CWE-Sys1 + CWE-UserF CWE-Sys2 + CWE-UserF
Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$
@@ II-F Procedure and Metrics
GPT-3.5 Turbo 0.45 0.14 0.12 0.12 0.38 0.14 0.12 0.12 0.52 0.34 0.35 0.32 0.44 0.24 0.23 0.21
GPT-4 Turbo 0.56 0.21 0.18 0.18 0.67 0.30 0.23 0.25 0.61 0.42 0.52 0.44 0.69 0.54 0.58 0.53
GPT-4o 0.62 0.23 0.19 0.20 0.52 0.22 0.19 0.19 0.67 0.61 0.69 0.61 0.65 0.60 0.74 0.63
CodeLlama-7b 0.22 0.12 0.09 0.08 0.29 0.08 0.05 0.06 0.11 0.13 0.10 0.08 0.13 0.11 0.08 0.06
CodeLlama-13b 0.33 0.09 0.08 0.08 0.25 0.08 0.05 0.05 0.23 0.09 0.07 0.07 0.21 0.10 0.09 0.07
Gemini 1.5 Pro 0.53 0.15 0.13 0.13 0.55 0.16 0.15 0.14 0.66 0.52 0.56 0.51 0.69 0.56 0.52 0.49
@@
GPT-3.5 Turbo 0.39 0.25 0.21 0.22 0.38 0.22 0.20 0.20 0.47 0.27 0.32 0.28 0.46 0.32 0.30 0.29
GPT-4 Turbo 0.61 0.43 0.42 0.40 0.66 0.42 0.35 0.36 0.51 0.53 0.54 0.48 0.59 0.59 0.55 0.52
GPT-4o 0.54 0.36 0.26 0.27 0.51 0.35 0.31 0.30 0.53 0.58 0.52 0.51 0.58 0.60 0.65 0.60
CodeLlama-7b 0.13 0.09 0.07 0.07 0.26 0.16 0.10 0.11 0.01 0.00 0.02 0.01 0.11 0.15 0.14 0.11
CodeLlama-13b 0.29 0.16 0.15 0.14 0.24 0.10 0.09 0.08 0.21 0.17 0.13 0.11 0.18 0.17 0.18 0.12
Gemini 1.5 Pro 0.54 0.24 0.21 0.20 0.50 0.25 0.22 0.22 0.67 0.61 0.64 0.58 0.71 0.72 0.62 0.62
@@
GPT-3.5 Turbo 0.43 0.21 0.20 0.19 0.34 0.20 0.19 0.19 0.44 0.46 0.43 0.39 0.41 0.32 0.33 0.31
GPT-4 Turbo 0.60 0.38 0.37 0.36 0.70 0.46 0.42 0.43 0.65 0.56 0.66 0.58 0.70 0.67 0.69 0.66
GPT-4o 0.61 0.37 0.31 0.32 0.48 0.33 0.28 0.29 0.66 0.64 0.74 0.65 0.60 0.62 0.76 0.64
CodeLlama-7b 0.22 0.19 0.15 0.15 0.19 0.08 0.06 0.07 0.16 0.12 0.20 0.13 0.15 0.12 0.17 0.12
CodeLlama-13b 0.31 0.18 0.15 0.15 0.22 0.15 0.12 0.12 0.22 0.14 0.13 0.12 0.24 0.20 0.21 0.18
Gemini 1.5 Pro 0.40 0.22 0.17 0.18 0.53 0.29 0.26 0.26 0.67 0.57 0.60 0.56 0.70 0.67 0.62 0.61
@@
GPT-3.5 Turbo 0.43 0.21 0.19 0.19 0.36 0.25 0.22 0.22 0.54 0.43 0.46 0.41 0.43 0.32 0.32 0.29
GPT-4 Turbo 0.53 0.39 0.39 0.37 0.62 0.44 0.42 0.41 0.64 0.59 0.68 0.60 0.68 0.57 0.66 0.57
GPT-4o 0.65 0.42 0.39 0.39 0.51 0.40 0.38 0.36 0.65 0.61 0.75 0.64 0.60 0.59 0.76 0.62
CodeLlama-7b 0.33 0.17 0.15 0.14 0.36 0.18 0.12 0.14 0.15 0.17 0.18 0.16 0.11 0.09 0.13 0.09
CodeLlama-13b 0.29 0.18 0.15 0.15 0.25 0.12 0.11 0.10 0.26 0.22 0.20 0.19 0.24 0.27 0.22 0.19
Gemini 1.5 Pro 0.58 0.32 0.28 0.29 0.62 0.32 0.30 0.30 0.60 0.45 0.40 0.39 0.64 0.45 0.42 0.41
@@
GPT-3.5 Turbo 0.53 0.25 0.26 0.24 0.47 0.27 0.27 0.25 0.64 0.42 0.50 0.44 0.50 0.41 0.40 0.37
GPT-4 Turbo 0.52 0.28 0.28 0.26 0.76 0.52 0.52 0.50 0.53 0.52 0.64 0.53 0.79 0.68 0.71 0.66
GPT-4o 0.69 0.37 0.42 0.38 0.53 0.29 0.29 0.27 0.74 0.59 0.62 0.59 0.69 0.53 0.61 0.55
CodeLlama-7b 0.21 0.16 0.20 0.17 0.36 0.25 0.25 0.22 0.14 0.13 0.19 0.13 0.14 0.05 0.12 0.06
CodeLlama-13b 0.41 0.26 0.30 0.26 0.38 0.27 0.27 0.23 0.17 0.08 0.07 0.06 0.14 0.13 0.11 0.09
Gemini 1.5 Pro 0.66 0.31 0.35 0.33 0.59 0.27 0.28 0.26 0.72 0.60 0.73 0.63 0.76 0.52 0.52 0.49
@@
GPT-3.5 Turbo 0.46 0.19 0.16 0.16 0.39 0.22 0.18 0.19 0.55 0.28 0.31 0.29 0.42 0.22 0.23 0.21
GPT-4 Turbo 0.51 0.29 0.23 0.25 0.63 0.42 0.35 0.37 0.68 0.49 0.55 0.50 0.69 0.48 0.52 0.48
GPT-4o 0.62 0.35 0.30 0.31 0.58 0.36 0.32 0.32 0.76 0.59 0.63 0.59 0.77 0.63 0.68 0.64
CodeLlama-7b 0.21 0.14 0.10 0.10 0.31 0.11 0.11 0.10 0.10 0.07 0.11 0.07 0.14 0.09 0.13 0.08
CodeLlama-13b 0.36 0.17 0.18 0.17 0.23 0.16 0.16 0.14 0.29 0.14 0.15 0.12 0.24 0.15 0.12 0.12
Gemini 1.5 Pro 0.51 0.28 0.26 0.26 0.51 0.29 0.26 0.26 0.65 0.59 0.69 0.59 0.65 0.52 0.60 0.52
III Experimental Results
III-A RQ1 (Vulnerability detection)
Table IV (vulnerability detection) presents the effectiveness results across the different prompt setups. Overall, GPT-4 Turbo and GPT-4o showed the best binary classification scores. The former achieved the highest accuracy and precision (0.81), whereas the latter showed the highest recall (1.00) and $F_{1}$ score (0.83). Changing the prompt from VD-Sys1 to VD-Sys2 (i.e., a more general role definition) had a positive impact only for GPT-3.5 Turbo and CodeLlama-13b, proving detrimental for the other LLMs. Although VD-Sys2 configurations achieved the highest recall results, they only marginally outperformed VD-Sys1 configurations and came at the cost of reduced precision. The best balance between detecting vulnerabilities and minimizing false positives was achieved by the prompt configurations VD-Sys1 + VD-User1 and VD-Sys1 + VD-User2 (the first two macro-columns). Among these, VD-Sys1 + VD-User1 was the most effective for minimizing false alarms (GPT-4 Turbo) and detecting security vulnerabilities (GPT-4o).
For Python, GPT-4 Turbo and GPT-4o consistently outperformed other models. In particular, both models achieved the highest $F_{1}$ score of 0.80 using the VD-User1 task prompt. Although Gemini 1.5 pro delivered the highest accuracy, its $F_{1}$ score did not reach more than 0.80 due to its lower recall than GPT-4 Turbo and GPT-4o across all the prompt settings.

### v2

@@ 2.3. Large Language Models
GPT-3.5 Turbo 0.59 0.77 0.67 0.55 0.82 0.66 0.73 0.70 0.71 0.58 0.77 0.66
GPT-4 Turbo 0.81 0.82 0.81 0.73 0.92 0.81 0.75 0.87 0.81 0.71 0.93 0.81
GPT-4o 0.72 0.97 0.83 0.64 0.99 0.78 0.57 1.00 0.72 0.57 0.99 0.73
CodeLlama-7b 0.58 0.89 0.70 0.54 0.93 0.68 0.52 0.98 0.68 0.55 0.84 0.66
CodeLlama-13b 0.76 0.28 0.41 0.67 0.65 0.66 0.62 0.77 0.69 0.62 0.81 0.71
Gemini 1.5 Pro 0.79 0.83 0.81 0.65 0.84 0.73 0.71 0.83 0.77 0.62 0.97 0.75
@@
GPT-3.5 Turbo 0.59 0.87 0.70 0.54 0.89 0.67 0.75 0.71 0.73 0.57 0.87 0.69
GPT-4 Turbo 0.73 0.79 0.76 0.72 0.89 0.80 0.73 0.87 0.80 0.66 0.87 0.75
GPT-4o 0.70 0.92 0.80 0.60 0.95 0.73 0.54 1.00 0.70 0.55 0.97 0.70
CodeLlama-7b 0.67 0.76 0.72 0.56 0.79 0.65 0.52 0.97 0.68 0.55 0.76 0.64
CodeLlama-13b 0.69 0.24 0.35 0.59 0.61 0.60 0.59 0.63 0.61 0.53 0.63 0.58
Gemini 1.5 Pro 0.79 0.71 0.75 0.59 0.71 0.64 0.70 0.84 0.76 0.60 0.95 0.73
@@
GPT-3.5 Turbo 0.55 0.68 0.61 0.48 0.70 0.57 0.68 0.73 0.70 0.56 0.68 0.61
GPT-4 Turbo 0.78 0.91 0.84 0.67 0.98 0.80 0.75 0.89 0.81 0.70 0.98 0.82
GPT-4o 0.75 1.00 0.85 0.63 1.00 0.77 0.58 1.00 0.73 0.59 1.00 0.75
CodeLlama-7b 0.55 0.93 0.69 0.55 0.98 0.70 0.51 0.95 0.67 0.51 0.86 0.64
CodeLlama-13b 0.93 0.32 0.47 0.67 0.77 0.72 0.60 0.86 0.71 0.72 0.89 0.80
Gemini 1.5 Pro 0.73 0.86 0.79 0.67 0.80 0.73 0.79 0.77 0.78 0.65 1.00 0.79
@@
GPT-3.5 Turbo 0.57 0.75 0.65 0.56 0.86 0.68 0.71 0.75 0.73 0.59 0.75 0.66
GPT-4 Turbo 0.82 0.89 0.85 0.76 0.97 0.85 0.71 0.94 0.81 0.73 0.97 0.83
GPT-4o 0.78 1.00 0.88 0.67 1.00 0.80 0.57 1.00 0.73 0.55 1.00 0.71
CodeLlama-7b 0.57 0.92 0.70 0.54 0.94 0.69 0.54 1.00 0.70 0.56 0.94 0.70
CodeLlama-13b 0.83 0.28 0.42 0.63 0.61 0.62 0.64 0.83 0.72 0.64 0.89 0.74
Gemini 1.5 Pro 0.85 0.92 0.88 0.67 0.89 0.76 0.67 0.89 0.76 0.61 0.97 0.75
@@
GPT-3.5 Turbo 0.69 0.86 0.77 0.63 0.93 0.75 0.85 0.76 0.80 0.67 0.90 0.76
GPT-4 Turbo 0.88 0.79 0.84 0.77 0.93 0.84 0.74 0.86 0.79 0.67 0.90 0.76
GPT-4o 0.61 0.97 0.75 0.59 1.00 0.74 0.52 1.00 0.68 0.52 0.97 0.67
CodeLlama-7b 0.58 0.97 0.73 0.53 0.97 0.68 0.51 1.00 0.67 0.60 0.90 0.72
CodeLlama-13b 0.57 0.28 0.37 0.72 0.62 0.67 0.64 0.72 0.68 0.58 0.86 0.69
Gemini 1.5 Pro 0.87 0.90 0.88 0.60 0.86 0.70 0.67 0.83 0.74 0.56 1.00 0.72
@@
GPT-3.5 Turbo 0.57 0.74 0.65 0.56 0.76 0.65 0.69 0.57 0.62 0.57 0.69 0.62
GPT-4 Turbo 0.86 0.71 0.78 0.76 0.83 0.80 0.83 0.81 0.82 0.78 0.93 0.85
GPT-4o 0.75 0.98 0.85 0.70 1.00 0.82 0.61 1.00 0.76 0.64 1.00 0.78
CodeLlama-7b 0.54 0.88 0.67 0.52 0.95 0.67 0.52 1.00 0.68 0.54 0.76 0.63
CodeLlama-13b 0.75 0.29 0.41 0.78 0.60 0.68 0.65 0.79 0.71 0.62 0.79 0.69
Gemini 1.5 Pro 0.74 0.76 0.75 0.70 0.93 0.80 0.74 0.83 0.79 0.65 0.95 0.77
2.4. Classification Methods
Our study examined three classification approaches.
Vulnerability Detection. A simple binary classification method in which the LLM is tasked to categorize an input snippet into either “vulnerable” or “not vulnerable”.
CWE classification (zero-shot). A multi-class classification method in which the LLM is tasked to categorize an input snippet as pertaining to a specific CWE. The zero-shot setting provides only a task description in natural language as part of the input prompt.
CWE classification (few-shot). Same multi-class classification method as before, but using a prompt that provides the top-25 CWE labels (1) to elicit CWE-specific reasoning and map the input snippet to the corresponding CWE category.
2.5. Prompt and Input Data
@@
In our study, the final prompts combine one system prompt with one user prompt. Table 4 reports all the prompts that were used for our experiments. We considered two system prompts: VD-Sys1 gives the LLM a detailed role for vulnerability detection. In contrast, VD-Sys2 had a more generic role, as described by Zhou et al. (Zhou et al., 2024b). Similarly, we designed two different types of user prompts to detect a vulnerability: U1 presents a concise request for vulnerability detection, whereas VD-User2 provides more descriptive instruction as used by Zhou et al. (Zhou et al., 2024b).
For the classification of CWE (RQ2), we introduced two system prompts (CWE-Sys1 and CWE-Sys2) and two user prompts (CWE-UserZ and CWE-UserF). CWE-Sys1 is a similar prompt compared to VD-Sys1, which provides a more specific role for CWE classification. Conversely, CWE-Sys2 is the same prompt as VD-Sys2 (Zhou et al., 2024b). CWE-UserZ is a user prompt for CWE classification with a zero-shot setting (i.e., without any examples). CWE-UserF is for CWE classification experiments with the few-shot setting, and it has a list of the top 25 CWEs comparable to CWE-UserZ. To form the final prompt text, we concatenated an input code snippet with one of our described prompt strategies.
2.6. Procedure and Metrics
To answer RQ1 and RQ2, we executed each LLM on all code snippets of our datasets, using the prompt strategies described in Section 2.5. As evaluation metrics, we employed four metrics, namely accuracy, recall, precision, and $F_{1}$ score, commonly used to evaluate the performance of LLMs in vulnerability detection. In our evaluation, valid detections/classifications were treated as the positive class, while detection/classification augmentations were considered the negative class. We accounted for CWE hierarchical relationships, considering positive scores for cases where the LLM predicted a code snippet labeled with a specific CWE (e.g., CWE-77) as one of its child CWEs (e.g., CWE-78), according to the MITRE CWE database (MITRE, 2024). For CWE classification (RQ2), we evaluated the overall performance by macro-averaging the precision parameter across all classes in the precision, recall, and $F_{1}$ score.
Overall, in our evaluation, we computed 18,144 predictions (6 LLMs $\times$ 378 code snippets $\times$ 8 prompt configurations), which accounted for 10 hours of computing time.
Table 6. RQ2: Effectiveness comparison of six LLMs using four prompt configurations for CWE classification (Precision, Recall, $F_{1}$).
LLM / Prompt C-Sys1 + UZ C-Sys2 + UZ C-Sys1 + UF C-Sys2 + UF
Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$
@@ 2.6. Procedure and Metrics
GPT-3.5 Turbo 0.14 0.12 0.12 0.14 0.12 0.12 0.34 0.35 0.32 0.24 0.23 0.21
GPT-4 Turbo 0.21 0.18 0.18 0.30 0.23 0.25 0.42 0.52 0.44 0.54 0.58 0.53
GPT-4o 0.23 0.19 0.20 0.22 0.19 0.19 0.61 0.69 0.61 0.60 0.74 0.63
CodeLlama-7b 0.12 0.09 0.08 0.08 0.05 0.06 0.13 0.10 0.08 0.11 0.08 0.06
CodeLlama-13b 0.09 0.08 0.08 0.08 0.05 0.05 0.09 0.07 0.07 0.10 0.09 0.07
Gemini 1.5 Pro 0.15 0.13 0.13 0.16 0.15 0.14 0.52 0.56 0.51 0.56 0.52 0.49
@@
GPT-3.5 Turbo 0.25 0.21 0.22 0.22 0.20 0.20 0.27 0.32 0.28 0.32 0.30 0.29
GPT-4 Turbo 0.43 0.42 0.40 0.42 0.35 0.36 0.53 0.54 0.48 0.59 0.55 0.52
GPT-4o 0.36 0.26 0.27 0.35 0.31 0.30 0.58 0.52 0.51 0.60 0.65 0.60
CodeLlama-7b 0.09 0.07 0.07 0.16 0.10 0.11 0.00 0.02 0.01 0.15 0.14 0.11
CodeLlama-13b 0.16 0.15 0.14 0.10 0.09 0.08 0.17 0.13 0.11 0.17 0.18 0.12
Gemini 1.5 Pro 0.24 0.21 0.20 0.25 0.22 0.22 0.61 0.64 0.58 0.72 0.62 0.62
@@
GPT-3.5 Turbo 0.21 0.20 0.19 0.20 0.19 0.19 0.46 0.43 0.39 0.32 0.33 0.31
GPT-4 Turbo 0.38 0.37 0.36 0.46 0.42 0.43 0.56 0.66 0.58 0.67 0.69 0.66
GPT-4o 0.37 0.31 0.32 0.33 0.28 0.29 0.64 0.74 0.65 0.62 0.76 0.64
CodeLlama-7b 0.19 0.15 0.15 0.08 0.06 0.07 0.12 0.20 0.13 0.12 0.17 0.12
CodeLlama-13b 0.18 0.15 0.15 0.15 0.12 0.12 0.14 0.13 0.12 0.20 0.21 0.18
Gemini 1.5 Pro 0.22 0.17 0.18 0.29 0.26 0.26 0.57 0.60 0.56 0.67 0.62 0.61
@@
GPT-3.5 Turbo 0.21 0.19 0.19 0.25 0.22 0.22 0.43 0.46 0.41 0.32 0.32 0.29
GPT-4 Turbo 0.39 0.39 0.37 0.44 0.42 0.41 0.59 0.68 0.60 0.57 0.66 0.57
GPT-4o 0.42 0.39 0.39 0.40 0.38 0.36 0.61 0.75 0.64 0.59 0.76 0.62
CodeLlama-7b 0.17 0.15 0.14 0.18 0.12 0.14 0.17 0.18 0.16 0.09 0.13 0.09
CodeLlama-13b 0.18 0.15 0.15 0.12 0.11 0.10 0.22 0.20 0.19 0.27 0.22 0.19
Gemini 1.5 Pro 0.32 0.28 0.29 0.32 0.30 0.30 0.45 0.40 0.39 0.45 0.42 0.41
@@
GPT-3.5 Turbo 0.25 0.26 0.24 0.27 0.27 0.25 0.42 0.50 0.44 0.41 0.40 0.37
GPT-4 Turbo 0.28 0.28 0.26 0.52 0.52 0.50 0.52 0.64 0.53 0.68 0.71 0.66
GPT-4o 0.37 0.42 0.38 0.29 0.29 0.27 0.59 0.62 0.59 0.53 0.61 0.55
CodeLlama-7b 0.16 0.20 0.17 0.25 0.25 0.22 0.13 0.19 0.13 0.05 0.12 0.06
CodeLlama-13b 0.26 0.30 0.26 0.27 0.27 0.23 0.08 0.07 0.06 0.13 0.11 0.09
Gemini 1.5 Pro 0.31 0.35 0.33 0.27 0.28 0.26 0.60 0.73 0.63 0.52 0.52 0.49
@@
GPT-3.5 Turbo 0.19 0.16 0.16 0.22 0.18 0.19 0.28 0.31 0.29 0.22 0.23 0.21
GPT-4 Turbo 0.29 0.23 0.25 0.42 0.35 0.37 0.49 0.55 0.50 0.48 0.52 0.48
GPT-4o 0.35 0.30 0.31 0.36 0.32 0.32 0.59 0.63 0.59 0.63 0.68 0.64
CodeLlama-7b 0.14 0.10 0.10 0.11 0.11 0.10 0.07 0.11 0.07 0.09 0.13 0.08
CodeLlama-13b 0.17 0.18 0.17 0.16 0.16 0.14 0.14 0.15 0.12 0.15 0.12 0.12
Gemini 1.5 Pro 0.28 0.26 0.26 0.29 0.26 0.26 0.59 0.69 0.59 0.52 0.60 0.52
3. Experimental Results
3.1. RQ1 (Vulnerability detection)
Table 5 (vulnerability detection) presents the effectiveness results across the different prompt setups. Overall, GPT-4 Turbo and GPT-4o showed the best binary classification scores. The former achieved the highest accuracy and precision (0.81), whereas the latter showed the highest recall (1.00) and $F_{1}$ score (0.83). Changing the prompt from Sys1 to VD-Sys2 (i.e., a more general role definition) had a positive impact only for GPT-3.5 Turbo and CodeLlama-13b, proving detrimental for the other LLMs. Although VD-Sys2 configurations achieved the highest recall results, they only marginally outperformed Sys1 configurations and came at the cost of reduced precision. The best balance between detecting vulnerabilities and minimizing false positives was achieved by the prompt configurations Sys1–U1 and Sys1 + VD-User2 (the first two macro-columns). Among these, Sys1–U1 was the most effective for minimizing false alarms (GPT-4 Turbo) and detecting vulnerabilities (GPT-4o).
For Python, GPT-4 Turbo and GPT-4o consistently outperformed other models. In particular, both models achieved the highest $F_{1}$ score of 0.80 using the U1 task prompt. Although Gemini 1.5 pro delivered the highest accuracy, its $F_{1}$ score did not reach more than 0.80 due to its lower recall than GPT-4 Turbo and GPT-4o.